Tiara Company has the following historical collection pattern for its credit sales: 70% collected in month of sale 15% collecPlease make sure the answer is correct 100% and to solve all parts clearly
Answer #1

a) Total Cash collection during October from October credit sales = 90000*70% = 63000

b) Total Cash collection during October from September credit sales = 80000*15% = 12000

c) Total Cash collections during October from August Credit sales = 70000*10% = 7000

d) Total cash collection during October from July Credit sales = 60000*4% = 2400

e) Total Cash collection = 2400+7000+12000+63000 = 84400